I am having trouble solving the following:
On the triangle, side A is x, side B is 16 and the hypotenuse is x+8. I can't seem to figure out where to start.

Answer by checkley71(8403)

DOE THE EQUATION A^2+B^2=C^2 LOOK FAMILIAR????
A=A SIDE
B=ANOTHER SIDE
C=THE HYPOTENUSE
HERE YOU HAVE A SIDE=X, SIDE 2=16 & THE HYPOTENUSE=X+8
X^2+18^2=(X+8)^2 NOW SOLVE FOR X
X^2+324=X^2+16X+64
X^2-X^2-16X=64-324
-16X=-260
X=260/-16
X=16.25
16.25+8=24.25 IS THE LENGTH OF THE HYPOTENUSE.
PROOF
16.25^2+18^2=24.25^2
264.0625+324=588.0625
588.0625=588.0625